# Interacting phantom dark energy: new accelerating scaling attractors

### v1

@@ I Introduction
Dark matter (DM) and dark energy (DE) are two main ingredients of the universe, constituting of 95% of its total energy budget. The DM sector is responsible for the observed structure formation, while the DE sector is in charge of the late time accelerated expansion Copeland et al. (2006); Bamba et al. (2012). The nature, origin and dynamics of both DE and DM are not yet clearly understood, nevertheless the $\Lambda$-Cold Dark Matter ($\Lambda$CDM) cosmological model, based on General Relativity, matches quite well with a large span of the astronomical datasets in which DE is a simple cosmological constant and DM is a cold pressureless sector. However, $\Lambda$CDM faces many challenges at both theoretical and observational levels Di Valentino et al. (2021a); Perivolaropoulos and Skara (2022); Abdalla et al. (2022), and among possible improvements is the consideration of a time varying nature in the DE sector.
@@
In general, DE is assumed to evolve independently and interact only with gravity. However, there is no fundamental reason to exclude possible interactions between DE and other fields, especially with the DM sector whose microscopic nature is also unknown. The interaction between DE and DM can alleviate the cosmic coincidence problem Amendola (2000); Chimento et al. (2003); Cai and Wang (2005); Pavon and Zimdahl (2005); Huey and Wandelt (2006); Hu and Ling (2006); del Campo et al. (2008, 2009), while it can potentially alleviate the current cosmological tensions too Di Valentino et al. (2021a); Perivolaropoulos and Skara (2022); Abdalla et al. (2022); Kumar and Nunes (2017); Di Valentino et al. (2017); Yang et al. (2018); Kumar et al. (2019); Pan et al. (2019a, b); Di Valentino et al. (2020a); Pan et al. (2020a); Pan and Yang (2023). Hence, over the last years interacting DM-DE scenarios have gained significant attention in the literature for their attractive features Guo et al. (2005); Sussman et al. (2005); Zimdahl (2005); Wang et al. (2005, 2006); Barrow and Clifton (2006); Setare (2006); Poplawski (2006); Sadjadi and Honardoost (2007); Chang et al. (2007); Kim et al. (2007); Rosenfeld (2007); Zimdahl and Pavon (2007); Feng et al. (2007); Cruz et al. (2008); Chen et al. (2008); Jamil et al. (2010); Jackson et al. (2009); Valiviita et al. (2010); Suwa and Nihei (2010); Karami and Ghaffari (2010); Wei (2010); Martinelli et al. (2010); Gavela et al. (2010); Lopez Honorez et al. (2010); Baldi (2011); Chen et al. (2014); Baldi and Salucci (2012); Chimento et al. (2013); Harko and Lobo (2013); Sun and Yue (2013); Li and Zhang (2014); Tamanini et al. (2014); Yang and Xu (2014); Li et al. (2014); Pan et al. (2015); Duniya et al. (2015); Li et al. (2016); Odderskov et al. (2016); Nunes et al. (2016); Ferreira et al. (2017); Biswas et al. (2017); Costa et al. (2017); Linton et al. (2018); Yang et al. (2019); Di Valentino et al. (2020b); Li et al. (2020); Yang et al. (2020); Pan et al. (2020b, c); Di Valentino et al. (2021b); Yang et al. (2021); Gao et al. (2021); Chatzidakis et al. (2022); Hou et al. (2023); Pan et al. (2023); Yang et al. (2023); Zhai et al. (2023); Bernui et al. (2023); Escamilla et al. (2023); Benisty et al. (2024); Halder et al. (2024a); Giarè et al. (2024a, b); Sabogal et al. (2024); Ghedini et al. (2024); Paliathanasis et al. (2024) (for reviews see Bolotin et al. (2014); Wang et al. (2016)).
@@
The interacting DM-DE models, widely known as Interacting DE (IDE) or Coupled DE, induce a richer dynamics in the universe evolution, recovering the non-interacting scenarios as special cases. The heart of the IDE models is the interaction function which characterizes the rate of energy and/or momentum transfer between the dark components, and consequently it affects the evolution of the universe. In order to investigate the dynamics of interacting cosmologies, one can confront them with observational datasets from various astronomical surveys and extract constraints on the involved functions and parameters. However, one can also apply the powerful approach of dynamical systems, which allows to extract analytical information independently of the initial conditions. The techniques of dynamical system analysis have been widely applied in the context of non-interacting and interacting dynamics Copeland et al. (1998); Billyard and Coley (2000); Gong et al. (2006); Chen and Gong (2009); Chen et al. (2009); Boehmer et al. (2008); Karwan (2008); Caldera-Cabral et al. (2009); Leon et al. (2009); Leon and Saridakis (2010); Xu et al. (2012); Leon and Saridakis (2013); Avelino et al. (2013); Boehmer et al. (2015a, b); Biswas and Chakraborty (2015a); Fadragas et al. (2014); Biswas and Chakraborty (2015b); Mahata and Chakraborty (2015); Banerjee and Roy (2015); Paliathanasis et al. (2015, 2016); Singh and Singh (2016); Dutta et al. (2017); Carneiro and Borges (2018); Odintsov et al. (2017); Zonunmawia et al. (2017); Bahamonde et al. (2018); Odintsov and Oikonomou (2018); Paliathanasis et al. (2019); Hernández-Almada et al. (2020); Sá (2020); Chakraborty et al. (2022); Paliathanasis et al. (2021); Sá (2021); Samart et al. (2021); Arevalo and Cid (2022); Sá (2024); Halder et al. (2024b).
@@
In the present work we are interested in performing a dynamical-system analysis of interacting cosmology considering a phantom dark energy Caldwell (2002); Dabrowski et al. (2003); Nojiri et al. (2005); Singh et al. (2003); Capozziello et al. (2006); Cai et al. (2010); Saridakis (2009). In particular, we consider various interacting phantom DE scenarios, characterized by four different potentials and two specific interacting functions, one corresponding to the global interaction rate and the other corresponding to the local interaction rate. The examination of new potentials even for the known and studied cases of interaction form reveals new accelerating scaling attractors that were never studied in the literature. This is an important result, since it was believed that in general, due to the phantom nature of the acceleration, accelerating scaling solutions in phantom cosmology were very rare.
@@ II Phantom Interacting Dark Energy
where $a(t)$ is the expansion scale factor of the universe. Concerning the gravitational theory we consider standard general relativity, and we additionally consider the standard (dark plus baryonic) matter sector of the universe corresponding to a perfect fluid. Concerning the dark energy sector, we attribute it to a phantom scalar field $\phi$ with action
@@
where a dot denotes derivative with respect to time. Note that the energy density of the phantom scalar field is unbounded from below and thus it may take negative values Bahamonde et al. (2018). Thus, the Friedmann equations are written as
@@
where $H=\dot{a}/a$ is the Hubble function (dots denote time-derivatives), $\kappa^{2}=8\pi G$ is the gravitational constant, and $\rho_{m}$, $p_{m}$ are respectively the energy density and pressure of the matter fluid.
@@
$\dot{\rho}_{m}+3H(p_{m}+\rho_{m})=-Q,$ (7)
$\dot{\rho}_{\phi}+3H(\rho_{\phi}+p_{\phi})=Q,$ (8)
where $Q$ is the interaction function which characterizes the transfer of energy between the dark sectors. For $Q>0$, the energy flow occurs from DM to DE, while for $Q<0$ energy flows in the reverse direction, namely from DE to DM. Notice that eqn. (8) can alternatively be expressed in a Klein-Gordon-like form, namely
@@
In this way, one can see that if the interaction function $Q(t)$ and the potential $V(\phi)$ of the scalar field are prescribed, then the evolution of the interacting fluids can be in principle determined. Hence, the choice of both of these functions is the starting point for any analysis.
In the following we focus on two interacting functions. The first interaction model is based on the most well-studied interaction form, namely Bolotin et al. (2014); Wang et al. (2016)
@@
where $\alpha$ is a dimensionless parameter. In this scenario the interaction rate is proportional to the energy density, which is expected to be the case since interactions in physics and chemistry are typically stronger if the number density of the involved particles increases. Nevertheless, the interaction rate is proportional to the Hubble function too, which although very convenient at the calculation stage raises the criticism that the DM-DE interaction rate, which is expected to depend on local physics, depends on a global feature of the whole Universe, namely the Hubble function Shafieloo et al. (2018). Hence, in this work we will additionally consider a second interaction function of the local form Shafieloo et al. (2018); Li et al. (2019)
@@
where $\Gamma$ is the dimensionfull coupling parameter. Finally, concerning the potential choices, we consider two forms that have been widely used in the literature, namely the well known exponential potential Copeland et al. (1998); Amendola (2000); Elizalde et al. (2004); Chen et al. (2009); Basilakos et al. (2011); Halder et al. (2024b) and the hyperbolic potential Paliathanasis et al. (2015), given by
@@
where $V_{0}>0$, and $\lambda$, $\eta$ are dimensionless free parameters. Another interesting potential form arises from the conformal transformation of Starobinsky $R^{2}$ gravity, and takes the form Ivanov et al. (2022)
@@
with $\mu$, $\beta$ dimensionless parameters. Finally, in order to cover the more general case $R+R^{2}+\Lambda$, we will also consider the potential Sebastiani et al. (2014)
@@
with $c_{0}$,$c_{1}$,$c_{2}$, the model parameters. Note that the last two potentials have not been studied in detail in the framework of phantom cosmology.
In summary, the background evolution of the scenario at hand is determined by the two Friedmann equations (5),(6), together with the conservation equations (7),(8) and the above choices for $Q$ and $V$. As usual, one can introduce the dark matter and dark energy density parameters as
@@
$w_{tot}$$\equiv$ $\frac{p_{\phi}+p_{m}}{\rho_{\phi}+\rho_{m}}=\frac{-\frac{1}{2}\dot{\phi}^{2}-V+w_{m}\rho_{m}}{-\frac{1}{2}\dot{\phi}^{2}+V+\rho_{m}}\,,$ (18)
with $w_{m}\equiv p_{m}/\rho_{m}$ the matter equation-of-state parameter.
@@ III Dynamical system analysis
In this section we will perform a detailed phase-space and stability analysis of interacting phantom cosmology. This approach allows one to obtain information on the global dynamics and the behavior of a cosmological scenario, independently of the initial conditions and the specific universe evolution Copeland et al. (1998). As usual, one first transforms the equations into an autonomous dynamical system, and extracts its critical points. Then, for each critical point one determines its stability by examining the sign of the eigenvalues of the perturbation matrix. The dynamical system analysis of interacting phantom cosmology has been performed in various articles Guo et al. (2005); Chen et al. (2009); Leon and Saridakis (2010); Shahalam et al. (2017); Chakraborty et al. (2022); Halder et al. (2024b), however it was restricted in specific potential choices. In this work we will perform it considering new potential forms and interaction models. In the following we focus on the case $w_{m}=0$, namely DM is pressureless and cold, and we examine the various potentials and interaction forms separately.
@@ III.1.1 Model I: $Q=\alpha H\rho_{m}$
  • The critical point $A_{2}$ exists in the parameter region shown in Table 1. Since the eigenvalues of the corresponding linearized matrix are complicated, the stability conditions have to be determined numerically, and are presented in Fig. 1. When $\alpha<-1$, it corresponds to an accelerating universe. Moreover, at $A_{2}$ we have $\Omega_{m}/\Omega_{\phi}\neq 0$, and thus this point represents late-time accelerating scaling solution. Hence, in this case, the coincidence problem can be alleviated, as reported earlier in Chen et al. (2009). Additionally, we mention that the critical points $A_{2}$ can represent purely matter dominated solutions (namely having $\Omega_{m}=1$) for the parameters $\lambda$ and $\alpha$ satisfying $\lambda^{2}=-(\alpha+3)^{2}/\alpha$. Thus, depending on the parameter values, point $A_{2}$ can correspond to the above possibilities. As we shall see below, for different potential choices such late-time stable accelerating scaling solutions can be obtained in a larger parameter region compared to this case.
@@ III.1.2 Model II: $Q=\Gamma\rho_{m}$
Finally, we mention that this local interaction model was investigated in Chen et al. (2009), considering the exponential potential, and the absence of accelerating scaling attractors was also reported.

### v2

@@ I Introduction
Dark matter (DM) and dark energy (DE) are two main ingredients of the universe, constituting of 95% of its total energy budget. The DM sector is responsible for the observed structure formation, while the DE sector is in charge of the late time accelerated expansion [1, 2]. The nature, origin and dynamics of both DE and DM are not yet clearly understood, nevertheless the $\Lambda$-Cold Dark Matter ($\Lambda$CDM) cosmological model, based on General Relativity, matches quite well with a large span of the astronomical datasets in which DE is a simple cosmological constant and DM is a cold pressureless sector. However, $\Lambda$CDM faces many challenges at both theoretical and observational levels [3, 4, 5], and among possible improvements is the consideration of a time varying nature in the DE sector.
@@
In general, DE is assumed to evolve independently and interact only with gravity. However, there is no fundamental reason to exclude possible interactions between DE and other fields, especially with the DM sector whose microscopic nature is also unknown. The interaction between DE and DM can alleviate the cosmic coincidence problem [6, 7, 8, 9, 10, 11, 12, 13], while it can potentially alleviate the current cosmological tensions too [3, 4, 5, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Hence, over the last years interacting DM-DE scenarios have gained significant attention in the literature for their attractive features [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91] (for reviews see [92, 93]).
@@
The interacting DM-DE models, widely known as Interacting DE (IDE) or Coupled DE, induce a richer dynamics in the universe evolution, recovering the non-interacting scenarios as special cases. The heart of the IDE models is the interaction function which characterizes the rate of energy and/or momentum transfer between the dark components, and consequently it affects the evolution of the universe. In order to investigate the dynamics of interacting cosmologies, one can confront them with observational datasets from various astronomical surveys and extract constraints on the involved functions and parameters. However, one can also apply the powerful approach of dynamical systems, which allows to extract analytical information independently of the initial conditions. The techniques of dynamical system analysis have been widely applied in the context of non-interacting and interacting dynamics [94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134].
@@
In the present work we are interested in performing a dynamical-system analysis of interacting cosmology considering a phantom dark energy [135, 136, 137, 138, 139, 140, 141, 142]. In particular, we consider various interacting phantom DE scenarios, characterized by four different potentials and two specific interacting functions, one corresponding to the global interaction rate and the other corresponding to the local interaction rate. The examination of new potentials even for the known and studied cases of interaction form reveals new accelerating scaling attractors that were never studied in the literature. This is an important result, since it was believed that in general, due to the phantom nature of the acceleration, accelerating scaling solutions in phantom cosmology were very rare.
@@ II Phantom Interacting Dark Energy
where $a(t)$ is the expansion scale factor of the universe. Concerning the gravitational theory we consider standard general relativity, and we additionally consider the DM sector of the universe corresponding to a perfect fluid. Concerning the dark energy sector, we attribute it to a phantom scalar field $\phi$ with action
@@
where a dot denotes derivative with respect to the cosmic time. Additionally, the Friedmann equations are written as
@@
where $H=\dot{a}/a$ is the Hubble function (dots denote time-derivatives) $\kappa^{2}=8\pi G$ is the gravitational constant, and $\rho_{m}$, $p_{m}$ are respectively the energy density and pressure of the DM fluid.
We mention here that in phantom cosmology the kinetic term is negative, however the sign of the total energy of the phantom field depends on its potential too. Hence, if one chooses potentials for which $V(\phi)>\dot{\phi}^{2}/2$ holds, then $\rho_{\phi}$ remains positive. However, even if the total phantom energy becomes negative, it does not lead to pathologies, since the matter energy density in (5) compensates this situation and thus $H^{2}$ remains always positive, as it should be. In other words, the fact that in phantom cosmology one can always impose a specific potential, and/or interaction form, that will make everything well-behaved and self-consistent (this is the case in many modified gravity theories, e.g. Horndeski gravity [143], where in principle they can lead to negative $H^{2}$, however one chooses the involved functions in a suitable way in order to always have $H^{2}>0$).
In summary, at the classical level the negativity of $\rho_{\phi}$ is not a problem, and all energy conditions are the same with the case of canonical fields. Nevertheless, issues might arise concerning the quantum field theory description of phantom fields, since unbounded-from-below states may lead to problems. However, in the literature there have been various attempts in constructing a phantom theory consistent with the basic requirements of quantum field theory, for example considering that the phantom fields arise as effective descriptions of more fundamental canonical theories [144, 145].
@@
$\dot{\rho}_{m}+3H(1+w_{m})\rho_{m}=-Q,$ (7)
$\dot{\rho}_{\phi}+3H(1+w_{\phi})\rho_{\phi}=Q,$ (8)
where $w_{m}=p_{m}/\rho_{m}$, $w_{\phi}=p_{\phi}/\rho_{\phi}$ are respectively the barotropic equation of state parameter of DM and DE, and $Q$ is the interaction function which characterizes the transfer of energy between the dark sectors. For $Q>0$, the energy flow occurs from DM to DE, while for $Q<0$ energy flows in the reverse direction, namely from DE to DM. Notice that eqn. (8) can alternatively be expressed in a Klein-Gordon-like form, namely
@@
In this way, one can see that if the interaction function $Q(t)$ and the potential $V(\phi)$ of the scalar field are prescribed, then the evolution of the interacting fluids can be determined in principle.
In summary, the choice of both functions $Q(t)$ and $V(\phi)$ is the starting point for any analysis. In particular, using the conservation equations, one can find the evolution laws of the dark fluids as
[…]
In the following we focus on two interacting functions. The first interaction model is based on the most well-studied interaction form, namely [92, 93]
@@
where $\alpha$ is a dimensionless parameter. In this scenario the interaction rate is proportional to the energy density of DM which is expected to be the case since interactions in physics and chemistry are typically stronger if the number density of the involved particles increases. Nevertheless, the interaction rate is proportional to the Hubble function too, which although very convenient at the calculation stage raises the criticism that the DM-DE interaction rate, which is expected to depend on local physics, depends on a global feature of the whole Universe, namely the Hubble function [146, 147, 148]. Hence, in this work we will additionally consider a second interaction function of the local form [146]
@@
where $\Gamma$ is the dimensionfull coupling parameter. Finally, concerning the potential choices, we consider two forms that have been widely used in the literature, namely the well known exponential potential [94, 6, 149, 98, 150, 134] and the hyperbolic potential [114], given by
@@
where $V_{0}>0$, and $\lambda$, $\eta$ are dimensionless free parameters. The third potential in this series is assumed to be [151]
@@
with $\mu$, $\beta$ dimensionless parameters. Let us mention here that such type of potential is not purely phenomenological, since similar forms may be derived in the context of $R^{2}$ gravity  [151]. Finally, we will also consider the following potential [152]
@@
with $c_{0}$,$c_{1}$,$c_{2}$, being the non-negative model parameters. Similarly to the previous potential, this potential choice can be inspired from $R+R^{2}+\Lambda$ gravity [152], and hence it may lead to interesting cosmological implications. We mention that although the above potentials have been inspired by higher-order theories of gravity, such as the $R^{2}$ one, in the present analysis we handle them as scalar-field potentials in the framework of General Relativity. Note that the last two potentials have not been studied in detail in the framework of phantom cosmology.
In summary, the background evolution of the scenario at hand is determined by the two Friedmann equations (5),(6), together with the conservation equations (7),(8) and the above choices for $Q$ and $V(\phi)$. As usual, one can introduce the dark matter and dark energy density parameters as
@@
$w_{tot}$$\equiv$ $\frac{p_{\phi}+p_{m}}{\rho_{\phi}+\rho_{m}}=\frac{-\frac{1}{2}\dot{\phi}^{2}-V(\phi)+w_{m}\rho_{m}}{-\frac{1}{2}\dot{\phi}^{2}+V(\phi)+\rho_{m}}\,.$ (20)
@@ III Dynamical system analysis
In this section we will perform a detailed phase-space and stability analysis of interacting phantom cosmology. This approach allows one to obtain information on the global dynamics and the behavior of a cosmological scenario, independently of the initial conditions and the specific universe evolution [94]. As usual, one first transforms the equations into an autonomous dynamical system, and extracts its critical points. Then, for each critical point one determines its stability by examining the sign of the eigenvalues of the perturbation matrix. The dynamical system analysis of interacting phantom cosmology has been performed in various articles [23, 98, 103, 153, 126, 134], however, it was restricted in specific potential choices. In this work we will perform it considering new potential forms and interaction models. In the following we focus on the case $w_{m}=0$, namely DM is cold, and we examine the various potentials and interaction forms separately.
@@ III.1.1 Model I: $Q=\alpha H\rho_{m}$
Let us make here the following comment. As the denominator in the evolution equation (25) contains $x$ explicitly, the system exhibits a singularity along the line $x=0$, which implies that the system becomes ill defined along the line $x=0$. Now, as $x=0$ marks a singular line, we can partition the domain $\mathbb{D}$ into two regions: one for $x>0$ and the other for $x<0$, denoted as $\mathbb{D}_{+}=\mathbb{D}\setminus\{x\leq 0\}$ and $\mathbb{D}_{-}=\mathbb{D}\setminus\{x\geq 0\}$, respectively. After applying a re-parametrization, as outlined in Appendix A.1, one can use blow-up techniques [154] to investigate the asymptotic dynamics of the trajectories near the singular line $x=0$ of the system (25)-(26) in the regions $\mathbb{D}_{+}$ and $\mathbb{D}_{-}$. Subsequently, the nature of the trajectories near the singular line $x=0$ of the system (25)-(26) can be characterized by blow-down. This itself is a subject for further investigation. Therefore, in the present work we exclude $x=0$ from our discussion.
@@
  • The critical point $A_{2}$ exists in the parameter region shown in Table 1. Since the eigenvalues of the corresponding linearized matrix are complicated, the stability conditions have to be determined numerically, and are presented in Fig. 1. When $\alpha<-1$, it corresponds to an accelerating universe. Moreover, at $A_{2}$ we have $\Omega_{m}/\Omega_{\phi}\neq 0$, and thus this point represents late-time accelerating scaling solution. Hence, in this case, the coincidence problem can be alleviated, as reported earlier in [98]. Additionally, we mention that the critical points $A_{2}$ can represent purely matter dominated solutions (namely having $\Omega_{m}=1$) for the parameters $\lambda$ and $\alpha$ satisfying $\lambda^{2}=-(\alpha+3)^{2}/\alpha$. Thus, depending on the parameter values, point $A_{2}$ can correspond to the above possibilities. As we shall see below, for different potential choices such late-time stable accelerating scaling solutions can be obtained in a larger parameter region compared to this case.
We close this subsection by mentioning that due to the non-compactness of the phase space in the $x-y$ direction discussed above, we should examine whether there are critical points at infinity. Hence, to complete the analysis, we compactify the phase space $\mathbb{D}$ by employing the Poincaré compactification technique [155, 154, 156, 121] and study the qualitative behaviour at the critical points at infinity. The details are provided in Appendix A.1. Our analysis reveals that the system possesses two critical points at infinity, namely $A^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0\right)$, which exhibit unstable behavior, and thus they cannot attract the Universe at late times. Additionally, we can see that at these critical points, $q\longrightarrow-\infty$. However, note that this is not considered as a Big Rip, since it is obtained at infinite rather than at finite time (see the classifications of singularities in a phantom Universe, in [138]).
@@ III.1.2 Model II: $Q=\Gamma\rho_{m}$
Similar to the previous subsection, here we also observe that the explicit dependence of the denominator in equation (29) on $x$ leads to a singularity at $x=0$ plane, rendering the above system ill-defined at this plane. This issue can be addressed in an asymptotic way by dividing the entire phase space $\mathbb{D}$ into two separate domains, namely, $\mathbb{D}_{+}=\mathbb{D}\setminus\{x\leq 0\}$ and $\mathbb{D}_{-}=\mathbb{D}\setminus\{x\geq 0\}$, and by using re-parametrization and suitable blow-up transformations. Similarly, equation (29) features $(1-z)$ in the denominator, resulting in a singularity at the $z=1$ plane, which makes the system ill-defined at this location too. To resolve this, the system can be regularized by introducing the factor $(1-z)$. However, the analysis of the regularized dynamical system and the technique utilized in this article are equivalent (see [99, 157, 121, 134, 84]), and therefore, the regularized system does not add any new physical implications in this context. Note that the same methodology can be applied to all autonomous systems in the following where the dynamical variables are in the denominators.
@@
Finally, since the phase space is not compact in the $x-y$ directions, we expect to have critical points at infinity. In order to provide a complete analysis we apply the Poincaré compactification technique to examine the stability at these critical points. This analysis shows that the system has two points at infinity, namely $B^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, both of which display unstable behavior, and thus they cannot attract the Universe at late times. The details are presented in Appendix A.2. At these points one can observe that $q\longrightarrow-\infty$.
Finally, we mention that this local interaction model was investigated in [98], considering the exponential potential, and the absence of accelerating scaling attractors was also reported.